# CoMoGaussian: Continuous Motion-Aware Gaussian Splatting from Motion-Blurred Images

### v1

@@ Abstract
3D Gaussian Splatting (3DGS) has gained significant attention for their high-quality novel view rendering, motivating research to address real-world challenges. A critical issue is the camera motion blur caused by movement during exposure, which hinders accurate 3D scene reconstruction. In this study, we propose CoMoGaussian, a Continuous Motion-Aware Gaussian Splatting that reconstructs precise 3D scenes from motion-blurred images while maintaining real-time rendering speed. Considering the complex motion patterns inherent in real-world camera movements, we predict continuous camera trajectories using neural ordinary differential equations (ODEs). To ensure accurate modeling, we employ rigid body transformations, preserving the shape and size of the object but rely on the discrete integration of sampled frames. To better approximate the continuous nature of motion blur, we introduce a continuous motion refinement (CMR) transformation that refines rigid transformations by incorporating additional learnable parameters. By revisiting fundamental camera theory and leveraging advanced neural ODE techniques, we achieve precise modeling of continuous camera trajectories, leading to improved reconstruction accuracy. Extensive experiments demonstrate state-of-the-art performance both quantitatively and qualitatively on benchmark datasets, which include a wide range of motion blur scenarios, from moderate to extreme blur.
@@ 4.1 CoMoGaussian Framework
Our goal is to reconstruct a sharp 3D scene using only camera motion-blurred images as input, thereby obtaining deblurred novel view images. Inspired by image blind deblurring methods, we follow the approach of Deblur-NeRF [20], learning a kernel that intentionally blurs images and excluding this kernel during rendering to produce sharp images. Since sharp image supervision is not available during training, we adhere to the fundamental principle of accurately simulating motion-blurred images. In this framework, learning to reconstruct blurry images inherently drives the model to capture the underlying sharp image representation. As illustrated in Fig. 1, our blurring process consists of camera poses along the camera motion trajectory, generated continuously in time order through neural ODEs [5]. Each pose is composed of a rigid body transformation $\mathbf{T}_{r}=[\mathbf{R}_{r}|\mathbf{t}_{r}]$ (Sec. 4.2), which maintains the shape and size of the subject, and a continuous motion refinement (CMR) transformation $\mathbf{T}_{c}=[\mathbf{R}_{c}|\mathbf{t}_{c}]$ (Sec. 4.3), which compensates for the limitations of approximating continuous motion using a discrete sum of rigid transformations by introducing learnable adjustments. Note that $\mathbf{R}\in\mathbb{R}^{3\times 3}$ and $\mathbf{t}\in\mathbb{R}^{3}$ represent rotation matrix and translation vector, and $\tilde{\mathbf{R}}\in\mathbb{R}^{3\times 3}$ and $\tilde{\mathbf{t}}\in\mathbb{R}^{3}$ stand for multiplicative offsets of change of the rotation matrix and translation vector, respectively. Given the two transformation matrices $\tilde{\mathbf{T}}_{r}=[\tilde{\mathbf{R}}_{r}|\tilde{\mathbf{t}}_{r}]$ and $\tilde{\mathbf{T}}_{c}=[\tilde{\mathbf{R}}_{c}|\tilde{\mathbf{t}}_{c}]$, the subsequent pose $\mathbf{T}^{\tau_{s}}(0\leq s<N)$ in the camera motion is derived as follows:
@@ 5.2 Ablation Study
To thoroughly demonstrate the validity and effectiveness of rigid body and CMR transformations, we conduct ablation experiments as shown in Tab. 3. Additionally, we perform further ablation studies for our continuous modeling approach, presented in Tab. 4. All the experiments for Tabs. 3 and 4 are conducted on Deblur-NeRF [20] real-world scenes. Additional ablation experiments are provided in the appendix.
@@
We analyze the impact of the main components of CoMoGaussian: the rigid body transformation and the CMR transformation. The baseline in Tab. 3 is Mip-Splatting [41], and when pixel-wise weights are not applied, the $N$ rendered images are averaged to produce a blurry image. We observe that incorporating the rigid body transformation alone into the baseline results in significantly improved performance across all metrics compared to the baseline. Furthermore, adding the CMR transformation to the rigid body motion, which compensates for the discretized integration in blurry pixel computation, consistently enhances performance across all metrics. Finally, applying the orthogonality loss $\mathcal{L}_{o}$, a component of the CMR transformation, leads to further improvements in LPIPS.
@@ 7 Conclusion
We propose CoMoGaussian, a novel method for reconstructing sharp 3D scenes from blurry images caused by camera motion. Under the assumption that objects remain static during camera movement, we apply a rigid body transformation and further enhance it with CMR transformation to compensate discretized numerical integration. These transformations are continuously modeled using neural ODEs, capturing continuous camera motion trajectories. Drawing inspiration from traditional image deblurring techniques and prior 3D scene deblurring research, we incorporate a pixel-wise weighting strategy with a lightweight CNN. CoMoGaussian surpasses state-of-the-art methods in 3D scene deblurring, with extensive experiments validating the effectiveness of each component.
@@ 8 Implementation Details
CoMoGaussian is trained for 40k iterations based on Mip-Splatting [41]. We set the number of poses $N$ that constitute the continuous camera trajectory to 9, which indicates that $M$ is 5. The embedding function of Sec. 4.2 is implemented by $\mathtt{nn.Embedding}$ of PyTorch, and the embedded features have the sizes of hidden state of 64. The sizes of hidden state of the single-layer encoders $\mathcal{E}_{r}$, $\mathcal{E}_{c}$, and the single-layer decoders $\mathcal{D}_{r}$, $\mathcal{D}_{c}$ are also 64. $\mathcal{D}_{r}$ consists of three head MLPs that extract the screw axis parameters $\omega$ and $v$, along with $\theta$, while $\mathcal{D}_{c}$ comprises two head MLPs that extract $\mathbf{A}_{c}$ and $\mathbf{t}_{c}$ as described in Sec. 4.3. The neural derivative function $f$ consists of two parallel single MLP layers, where one is designated for the rotation component and the other for the translation component, where $f$ and $g$ share the learnable parameters. To ensure nonlinearity in the camera motion within the latent space, we apply the $\mathtt{ReLU}$ activation function to each layer. The CNN $\mathcal{F}$ consists of two convolutional layers with 64 channels with kernel size of 5$\times$5 for the first layer and 3$\times$3 for the rest one. The pixel-wise weights are obtained by applying a pointwise convolutional layer to the output of $\mathcal{F}$, and the scalar mask $\mathcal{M}$ is obtained by averaging it to the batch axis and applying another pointwise convolution to the output of $\mathcal{F}$. For first 1k iterations, Gaussian primitives are roughly trained without rigid body transformation and CMR transformation. After 1k iterations, those transformations start to be trained without the pixel-wise weight and the scalar mask to allow the initial camera motion path to be sufficiently optimized. After 3k iterations, the pixel-wise weight and the scalar mask start training. We set $\lambda_{c}$, $\lambda_{o}$, and $\lambda_{\mathcal{M}}$ to 0.3, $10^{-4}$, and $10^{-3}$ respectively for the objective function. All experiments are conducted on a single NVIDIA RTX 4090 GPU.
@@ 13 Derivation of Rigid Body Motion [19]
$[\hat{\omega}]=\left[\begin{array}[]{ccc}0&-\hat{\omega}_{z}&\hat{\omega}_{y}\\ \hat{\omega}_{z}&0&-\hat{\omega}_{x}\\ -\hat{\omega}_{y}&\hat{\omega}_{x}&0\\ \end{array}\right]\in\mathfrak{so}(3),~{}where~{}[\hat{\omega}]^{3}=I.$ (17)

### v2

@@ Abstract
3D Gaussian Splatting (3DGS) has gained significant attention due to its high-quality novel view rendering, motivating research to address real-world challenges. A critical issue is the camera motion blur caused by movement during exposure, which hinders accurate 3D scene reconstruction. In this study, we propose CoMoGaussian, a Continuous Motion-Aware Gaussian Splatting that reconstructs precise 3D scenes from motion-blurred images while maintaining real-time rendering speed. Considering the complex motion patterns inherent in real-world camera movements, we predict continuous camera trajectories using neural ordinary differential equations (ODEs). To ensure accurate modeling, we employ rigid body transformations, preserving the shape and size of the object but rely on the discrete integration of sampled frames. To better approximate the continuous nature of motion blur, we introduce a continuous motion refinement (CMR) transformation that refines rigid transformations by incorporating additional learnable parameters. By revisiting fundamental camera theory and leveraging advanced neural ODE techniques, we achieve precise modeling of continuous camera trajectories, leading to improved reconstruction accuracy. Extensive experiments demonstrate state-of-the-art performance both quantitatively and qualitatively on benchmark datasets, which include a wide range of motion blur scenarios, from moderate to extreme blur. Project page is available at https://Jho-Yonsei.github.io/CoMoGaussian.
@@ 4.1 CoMoGaussian Framework
Our goal is to reconstruct a sharp 3D scene using only camera motion-blurred images as input, thereby obtaining deblurred novel view images. Inspired by image blind deblurring methods, we follow the approach of Deblur-NeRF [20], learning a kernel that intentionally blurs images and excluding this kernel during rendering to produce sharp images. Since sharp image supervision is not available during training, we adhere to the fundamental principle of accurately simulating motion-blurred images. In this framework, learning to reconstruct blurry images inherently drives the model to capture the underlying sharp image representation. As illustrated in Fig. 1, our blurring process consists of camera poses along the camera motion trajectory, generated continuously in time order through neural ODEs [5]. Each pose is composed of a rigid body transformation $\tilde{\mathbf{T}}_{r}=[\tilde{\mathbf{R}}_{r}|\tilde{\mathbf{t}}_{r}]$ (Sec. 4.2), which maintains the shape and size of the subject, and a continuous motion refinement (CMR) transformation $\tilde{\mathbf{T}}_{c}=[\tilde{\mathbf{R}}_{c}|\tilde{\mathbf{t}}_{c}]$ (Sec. 4.3), which compensates for the limitations of approximating continuous motion using a discrete sum of rigid transformations by introducing learnable adjustments. Note that $\mathbf{R}\in\mathbb{R}^{3\times 3}$ and $\mathbf{t}\in\mathbb{R}^{3}$ represent rotation matrix and translation vector, and $\tilde{\mathbf{R}}\in\mathbb{R}^{3\times 3}$ and $\tilde{\mathbf{t}}\in\mathbb{R}^{3}$ stand for multiplicative offsets of change of the rotation matrix and translation vector, respectively. Given the two transformation matrices $\tilde{\mathbf{T}}_{r}=[\tilde{\mathbf{R}}_{r}|\tilde{\mathbf{t}}_{r}]$ and $\tilde{\mathbf{T}}_{c}=[\tilde{\mathbf{R}}_{c}|\tilde{\mathbf{t}}_{c}]$, the subsequent pose $\mathbf{T}^{\tau_{s}}(0\leq s<N)$ in the camera motion is derived as follows:
@@ 5.2 Ablation Study
To thoroughly demonstrate the effectiveness of rigid body and CMR transformations, we conduct ablation experiments as shown in Tab. 3. Additionally, we perform further ablation studies for our continuous modeling, presented in Tab. 4. All the experiments for Tabs. 3 and 4 are conducted on Deblur-NeRF [20] real-world scenes. Additional ablation experiments are provided in the appendix.
@@
We analyze the impact of the main components of CoMoGaussian: the rigid body transformation and the CMR transformation. The baseline in Tab. 3 is Mip-Splatting [41], and when pixel-wise weights are not applied, the $N$ rendered images are averaged to produce a blurry image. We observe that incorporating the rigid body transformation alone into the baseline results in significantly improved performance across all metrics compared to the baseline. Furthermore, adding the CMR transformation to the rigid body motion, which compensates for the discretized integration, consistently enhances performance across all metrics. Finally, applying the orthogonality loss $\mathcal{L}_{o}$, a component of the CMR transformation, leads to further improvements in LPIPS.
@@ 7 Conclusion
We propose CoMoGaussian, a novel method for reconstructing sharp 3D scenes from blurry images caused by camera motion. We apply a rigid body transformation and further enhance it with CMR transformation to compensate discretized numerical integration. These transformations are continuously modeled using neural ODEs, capturing continuous camera motion trajectories. Drawing inspiration from prior 3D scene deblurring research, we incorporate a pixel-wise weighting strategy with a lightweight CNN. CoMoGaussian surpasses state-of-the-art methods in 3D scene deblurring, with extensive experiments validating the effectiveness of each component.
Acknowledgements.
This work was supported by the Yonsei Signature Research Cluster Program of 2024 (2024-22-0161), the National Research Foundation of Korea (NRF) grant funded by the Korean government (MSIT)(No. RS-2024-00340745), and the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT)(No. RS-2024-00423362).
@@ 8 Implementation Details
CoMoGaussian is trained for 40k iterations based on Mip-Splatting [41]. We set the number of poses $N$ that constitute the continuous camera trajectory to 9, which indicates that $M$ is 5. The embedding function of Sec. 4.2 is implemented by $\mathtt{nn.Embedding}$ of PyTorch, and the embedded features have the sizes of hidden state of 64. The sizes of hidden state of the single-layer encoders $\mathcal{E}_{r}$, $\mathcal{E}_{c}$, and the single-layer decoders $\mathcal{D}_{r}$, $\mathcal{D}_{c}$ are also 64. $\mathcal{D}_{r}$ consists of three head MLPs that extract the screw axis parameters $\omega$ and $v$, along with $\theta$, while $\mathcal{D}_{c}$ comprises two head MLPs that extract $\mathbf{A}_{c}$ and $\mathbf{t}_{c}$ as described in Sec. 4.3. The neural derivative function $f$ consists of two parallel single MLP layers, where one is designated for the rotation component and the other for the translation component, where $f$ and $g$ share the learnable parameters. To ensure nonlinearity in the camera motion within the latent space, we apply the $\mathtt{ReLU}$ activation function to each layer. The CNN $\mathcal{F}$ consists of two convolutional layers with 32 channels with kernel size of 5$\times$5 for the first layer and 3$\times$3 for the rest one. The pixel-wise weights are obtained by applying a pointwise convolutional layer to the output of $\mathcal{F}$, and the scalar mask $\mathcal{M}$ is obtained by averaging it to the batch axis and applying another pointwise convolution to the output of $\mathcal{F}$. For first 1k iterations, Gaussian primitives are roughly trained without rigid body transformation and CMR transformation. After 1k iterations, those transformations start to be trained without the pixel-wise weight and the scalar mask to allow the initial camera motion path to be sufficiently optimized. After 3k iterations, the pixel-wise weight and the scalar mask start training. We set $\lambda_{c}$, $\lambda_{o}$, and $\lambda_{\mathcal{M}}$ to 0.3, $10^{-4}$, and $10^{-3}$ respectively for the objective function. All experiments are conducted on a single NVIDIA RTX 4090 GPU.
@@ 13 Derivation of Rigid Body Motion [19]
$[\hat{\omega}]=\left[\begin{array}[]{ccc}0&-\hat{\omega}_{z}&\hat{\omega}_{y}\\ \hat{\omega}_{z}&0&-\hat{\omega}_{x}\\ -\hat{\omega}_{y}&\hat{\omega}_{x}&0\\ \end{array}\right]\in\mathfrak{so}(3),~{}where~{}[\hat{\omega}]^{3}=-[\hat{\omega}].$ (17)